OpenStudy (sophtlam):

2 six sided dice are rolled. the rolls re independent. what is the probability that at least one die rolls a 5?

OpenStudy (sophtlam):

are*

OpenStudy (sophtlam):

the only way i can do this question is by drawing out a table, but id like to be able to understand a formula

OpenStudy (sophtlam):

answer is 11/36

OpenStudy (mathmate):

 The outcomes are marked, as you have probably done. The number eleven / 36 is obtained by adding 1. probability of having once die rolling a 5 (1/6) 2. probability of the other die rolling a 5 (1/6) LESS 3. probability of both dice rolling a 5 (A and B) = 1/6*1/6=1/36 We need to subtract P(A&B) because we have counted it twice in items 1 & 2. You can see this from the diagram: horizontal and vertical rows intersect. That's the one that is counted twice. In general, it is: P(A OR B) = P(A)+P(B)-P(A AND B)